house A is 2years older than house B. nine years ago house A was 3 times as old as house B. how old is house A and house B.

Answer by ptaylor(2198)

Let x=B's current age; x-9=B's age 9 years ago
Then x+2=A's current age; x+2-9=x-7=A's age 9 years ago
And we are told that:
x-7=3(x-9)
x-7=3x-27
2x=20
x=10 --------------B's current age
x+2=10+2=12 A's current age
CK
12-9=3*(10-9)
3=3
hope this helps---ptaylor